OpenStudy (anonymous):

The difference of two numbers is 3. Their sum is 13. Find the numbers

OpenStudy (anonymous):

guess and check try 7 and 6 no how about 5 and 8?

OpenStudy (anonymous):

ty

myininaya (myininaya):

difference of two numbers is 3 x-y=3 x-y is the difference of x and y and the difference is 3 the sum of the two numbers is 13 x+y=13

myininaya (myininaya):

we have two equations x-y=3 and x+y=13

myininaya (myininaya):

add together x-y=3 x+y=13 ------- 2x+0y=16 2x=16 x=8 now plug x=8 into one of the equations we got above and solve for y

OpenStudy (anonymous):

8 and 5